Hausmitteilung

An alle Mitarbeiterinnen und Mitarbeiter Organisationseinheit

(zweizeilig)

Position Name

Telefon (03 51) 2 60-00 00 Telefax (03 51) 2 60-00 00 V.Name@fz-rossendorf.de

16. Januar 2001

Sehr geehrte ...

Text

signname